### v2

# Machine-learned nodal structures of Fermion systems

Abstract
Simulating strongly correlated fermionic systems remains a fundamental challenge in quantum physics, largely due to the sign problem in quantum Monte Carlo (QMC) methods. We present a neural network-based variational Monte Carlo (NN-VMC) approach, leveraging a flexible neural network ansatz to represent the many-body wavefunction. Focusing on quantum dots with up to 30 electrons, we demonstrate that NN-VMC significantly reduces variational bias and achieves ground-state energies surpassing those of fixed-node diffusion Monte Carlo (DMC). A key feature is that the neural network adaptively learns and optimizes nodal structures during energy minimization. We provide qualitative insights into the nodal structure of fermionic wavefunctions by comparing the nodal structures generated by NN-VMC with those obtained from traditional trial functions. Additionally, we reveal spin-resolved radial distributions and electron density profiles, highlighting the versatility and accuracy of NN-VMC. This work underscores the potential of machine learning to advance quantum simulations and deepen our understanding of strongly correlated systems.
Introduction - Accurately modeling strongly correlated fermionic systems remains a central challenge in quantum many-body physics, with implications for systems ranging from quantum chromodynamics and nuclear matter to ultracold gases and condensed matter systems [1, 2, 3, 4, 5]. These systems exhibit strong correlations that traditional methods struggle to capture, requiring many-body correlations in the trial function. Quantum Monte Carlo (QMC) methods [6] have proven particularly effective, as they leverage non-relativistic quantum mechanics to capture intricate many-body effects.
Quantum dots (QDs) are versatile platforms for studying strong correlations and excitonic states [7, 8, 9], with applications in scalable quantum computing and other quantum technologies [10, 11, 12, 13, 14, 15]. Furthermore, QDs provide a unique platform for studying generalized Fermi-Hubbard physics, offering insights into condensed matter phenomena and advancing our understanding of quantum many-body systems [16, 17].
Machine learning (ML) offers a powerful alternative to traditional trial functions in QMC methods by enabling more flexible and expressive wavefunction representations. Deep neural networks (DNNs) capture complex many-body correlations more effectively than conventional approaches, enhancing wavefunction accuracy and reducing variational bias [18, 19, 20]. This adaptability makes ML-based methods particularly valuable for refining nodal structures and mitigating the limitations of fixed-node approximations, which constrain random walks to regions where a predefined wavefunction retains a single sign. By learning more accurate nodal surfaces, ML approaches achieve better approximations of the true wavefunction [21, 22], ultimately improving the accuracy of QMC simulations.
The fixed-node approximation mitigates the sign problem but introduces an intrinsic bias due to the predetermined nodal structure [23, 24, 25, 26, 27]. Its accuracy depends heavily on the quality of the trial wavefunction´s nodal surfaces, as highlighted by Ceperley [28]. The computational complexity of the sign problem has been widely studied, with Troyer and Wiese [29] analyzing its fundamental limitations. Additionally, Foulkes et al. [30] examined its impact on solid-state simulations, providing further context for its challenges. Moreover, traditional trial wavefunctions impose rigid nodal surfaces, limiting their flexibility and accuracy. This is particularly evident in regions near the nodal hypersurfaces, where traditional trial functions often exhibit significant inaccuracies.
Building on these insights, Carleo and Troyer [31] revolutionized the field by introducing neural network quantum states (NNQS), demonstrating their capacity to accurately represent many-body wavefunctions and solve quantum spin systems. This breakthrough paved the way for applications such as deep neural network solutions for the electronic Schrödinger equation [32] and the extension [21] of these methods to ab initio simulations of many-electron systems. Furthermore, Luo and Clark [33] demonstrated the power of neural networks in refining nodal structures through backflow transformations, while Choo et al. [34] explored how NNQS can capture symmetries and excitations in many-body systems. These contributions highlight the transformative potential of ML-based approaches in overcoming the limitations of traditional QMC methods, particularly in addressing the sign problem and enhancing the accuracy of quantum many-body simulations.
In this work, we employ neural network-based variational Monte Carlo (NN-VMC) [21] to implicitly optimize nodal structures during energy minimization. Focusing on quantum dot systems with up to 30 electrons, we demonstrate that NN-VMC reduces variational bias and achieves ground-state energy estimates that surpass those obtained with diffusion Monte Carlo (DMC) using the fixed-node approximation. More importantly, we provide qualitative insights into the nodal structure of fermionic wavefunctions. Unlike previous works on molecular systems, which did not compare the nodal structures generated by NN-VMC and traditional trial functions [22, 35], our study provides such a comparison. Furthermore, we provide new insights into strongly correlated quantum dot systems, such as spin-resolved radial distribution functions and electron density profiles, highlighting the broader applicability of ML-based wavefunction representations in quantum many-body physics. Our findings establish the broader applicability and effectiveness of NN-VMC methods for studying strongly correlated fermionic systems, marking a significant step forward in the field. Moreover, these NN representations can be efficiently computed using graphical process units (GPUs), and their derivatives can be calculated without finite difference errors using automatic differentiation.
Given the challenges associated with QMC methods and the crucial role of nodal structure optimization, we adopt a neural network-based variational Monte Carlo (NN-VMC) approach. By leveraging the flexibility of deep neural networks, we refine wavefunction representations to improve accuracy in strongly correlated fermionic systems. In the next section, we detail the NN-VMC framework, including the modeling of trial wavefunctions, optimization strategies, and key computational techniques applied to quantum dot systems.
Methods - Our NN-VMC method models trial wavefunctions using a neural network ansatz, optimizing them based on variational principles and computing physical properties through Monte Carlo sampling. Below, we describe the system model, Hamiltonian, neural network architecture, and the optimization process that enables adaptive nodal structure refinement.
We investigate a system of circular QDs made from semiconductor heterostructures, which is modeled by electrons interacting via the Coulomb potential and confined within a two-dimensional harmonic trap. With electrons placed in the plane, this model’s Hamiltonian is
where is the effective mass the electron, is the reduced Planck constant, is the angular frequency of the harmonic trap, is the dielectric constant of the semiconductor, is the electron charge. The indices and identify two spatial points corresponding to the electrons in the system. The distance between these points is given by where . It is convenient to define as the set of all such points. A dimensionless Hamiltonian form is used in this work is obtained by the introduction of an effective Hartree energy scale , where is the effective Bohr radius, and is the Coulomb constant, see Section I of the Supplemental Material (SM).
We choose parameters based on a quantum dot of gallium arsenide, GaAs, which has an electronic effective mass ( is the electron mass), and a dielectric constant [36]. For the confining trap, we chose , a typical value in the literature investigating the behavior of quantum dots [37].
To represent the wave function of this fermionic system, we employ a neural network architecture based on the FermiNet approach [38, 39, 40], which provides a highly flexible representation of complex wave functions. Despite requiring extensive training, this approach offers superior representational power compared to other leading neural network architectures for fermionic systems [32, 33]. The FermiNet-based trial function was then used to perform QMC variational simulations, coupled with iterative optimization steps. The trial wave function is given by
where is a variational parameter and orbitals are constructed using artificial NNs. The indices and identify spatial points for electrons with either spin-up or spin-down, where . The neural network output is associated with the spatial point that corresponds to the same spin of , while are variational parameters.
where is not associate to any spin and the spin of index is the same of .
The Kronecker-factored approximate curvature (KFAC) algorithm [41] is used to optimize the variational parameters of the trial wavefunction, ensuring stable convergence.
A schematic diagram of the NN-VMC method is included in the SM. It is important to note that for , the harmonic confinement potential used in this study does not introduce essential singularities. The inclusion of in the ansatz is designed to enhance the flexibility of the neural network model to adapt to different spatial configurations rather than to address any singularity in the trap potential. In two and three dimensions, the kinetic energy operator inherently includes terms proportional to , which naturally regulate divergences at small . An additional dependence in the kinetic energy arising from the trial function does not introduce unbounded states in this case. The neural network-based approach ensures smooth approximations of the wavefunction through activation functions and optimization mechanisms that minimize energy while satisfying the variational principle.
Any potential discontinuity at is suppressed during training, as contributions of non-physical kinetic energy are penalized in the optimization process. The variational parameters, including those that influence terms involving , are adjusted to minimize energy, ensuring that the trial function remains physically valid. The resulting energies and statistical uncertainties confirm that the ansatz does not lead to unbounded kinetic energy contributions. However, further analysis of the behavior of terms near could provide deeper theoretical insights into their impact on wavefunction optimization.
The trial wave function is parameterized using a neural network ansatz following the FermiNet architecture, as given in Eqs. (2)–(3). The input features, layer propagation equations, and interaction terms that define the one- and two-electron streams are detailed in the SM. These include the transformation of electron coordinates, activation functions, and network layer updates necessary for efficient energy optimization.
This methodology enables the adaptive optimization of nodal structures and the accurate computation of ground state energies, as demonstrated in the results section.
Results - We investigated the effectiveness of using machine learning, specifically neural networks, in representing wave functions for this fermionic system. Our focus was on QDs with up to 30 electrons, and we compared the results of our neural network-based variational Monte Carlo (NN-VMC) approach with the diffusion Monte Carlo (DMC) method using the fixed-node approximation.
The evolution of the total variational energy for a 12-electron QD is shown in Fig. 1. The energy decreases rapidly initially, followed by a slower decline, reaching values below DMC results after about 40,000 optimization steps. Further refinement requires hundreds of thousands of steps, highlighting the computational cost of high accuracy. As the electron count increases, the computational burden grows significantly, making calculations for systems with more than 30 electrons demanding. Nevertheless, NN-VMC achieves a statistically significant improvement (0.5%) over DMC, with smaller energy uncertainties. This improvement stems from NN-VMC’s ability to adaptively refine nodal surfaces, capturing complex correlations that traditional methods struggle to represent. Compared to PIMC, NN-VMC also reduces kinetic energy contributions (see SM).
The hyperparameters were carefully selected based on empirical tuning. Their typical values are presented in Section IV of the SM, along with a sensitivity analysis of their impact on the results.
After the optimization process, we select the optimized set of variational parameters and perform a standard variational Monte Carlo integration to further reduce statistical uncertainties. Table I compares the total energies obtained by NN-VMC with those obtained by the leading methods used in the investigation of QDs, DMC, path-integral Monte Carlo (PIMC), coupled cluster with singles and doubles (CCSD) and in-medium similarity renormalization group (IM-SRG). Particularly, the CCSD and IM-SRG methods do not provide an upper bound for the total energy of the system. Across all electron counts, our approach consistently yields lower energy values, except when comparing with IM-SRG for a few electrons. For the 30-electron system, the total energy calculated using our neural network approach is 123.9469(2) , compared to 123.9683(2) from DMC, representing an improvement of 0.02 . Our results highlight the consistent improvements achieved by the NN-VMC method.
This work | 7.59703(9) | 25.62599(6) | 61.9073(1) | 123.9372(1) |
DMC111[42], ∗[43]; b[44]; c[45] | 7.6001(1) | 25.6356(1) | 61.922(2) | 123.9683(2)∗ |
PIMCb | 7.5980(1) | 25.6456(6) | 61.992(3) | N/A |
CCSDc | 7.6341 | 25.7345 | 62.1312 | 124.3630 |
IM-SRGc | 7.5731 | 25.6259 | 61.9585 | 124.1041 |
To evaluate the nodal structures generated by our neural network representation, we compared the wave function with a harmonic oscillator trial function, the usual Slater determinant of non-interacting harmonic oscillator orbitals with the harmonic frequency as a variational parameter. The primary objective of this study is to demonstrate the effectiveness of NN-VMC in reducing variational bias and achieving lower ground state energies compared to traditional methods like DMC. The energy minimization process is the central metric for evaluating the success of your approach. A detailed quantitative analysis of the nodal structure, while valuable, is not directly necessary to achieve this primary goal. Moreover, the sophisticated analytical tools necessary for such quantitative analysis would significantly extend the scope of the study. The qualitative comparison of nodal structures provided in the study offers sufficient initial validation of the NN-VMC method’s ability to produce smoother and more symmetric nodal surfaces. These insights highlight the adaptability of neural networks in reducing variational bias.
The nodal structure is not explicitly prescribed but is implicitly learned by the neural network through energy minimization. Defining a generalizable quantitative curvature metric is non-trivial and remains an open area of research [35]. As the network adjusts its parameters during training, it optimizes the placement and smoothness of the nodal surfaces to reduce the variational energy, leading to a self-consistent nodal topology that reflects the underlying physics of the system, Fig. 2 illustrates the smooth and symmetric nodal surfaces generated by NN-VMC. In contrast, traditional trial functions impose rigid nodal structures that lead to higher energy estimates (see the SM).
The nodal regions of the neural network-derived hyperplanes in blue (right) are smoother and exhibit less curvature compared to the traditional trial function represented in green (left), leading to a reduction in kinetic energy and an overall more accurate representation of the system. The comparison was repeated by choosing different initial electronic configurations, also including electrons with spin down. The result was always consistent with the one we show in Fig. 2. Naturally, changing the electron being moved impacts the overall aspect of the corresponding nodal structure, as can be seen in Section V of the SM. The NN representation of the trial function shows the requirements of a “good” nodal structure, i.e. it exhibits smoothness and symmetry. Traditional trial wavefunctions often impose rigid nodal surfaces, which can lead to high kinetic energy contributions and suboptimal energy estimates. In contrast, the neural network’s flexibility allows it to adapt the nodal surfaces smoothly and symmetrically, as qualitatively depicted in Fig. 2.
Showcasing the ability of the neural-network wave function to yield accurate results for physical properties of the system, in Fig. 3 we show results for radial distribution functions of different relative electronic spin orientations. Furthermore, the two dimensional density profile for a QD can be computed through
normalized by imposing . The result for a QD with 20 electrons is displayed in Fig. 4. Although circular symmetry was not explicitly imposed, our neural network was able to capture this feature during training. This reinforces the flexibility and effectiveness of our approach in representing complex wave functions without prior assumptions about the geometry of the system. Figs. 1-4 provide new insights into the interplay between confinement potentials and electron correlations, showcasing the versatility of NN-VMC.
The small energy uncertainties (e.g., 0.00009 for a 12-electron system) highlight the robustness of the NN-VMC method and confirm the statistical significance of its improvements over DMC. These results confirm that our machine learning-based approach can reliably reduce variational bias in quantum Monte Carlo simulations.
Conclusion - Several sophisticated approaches have been proposed to mitigate or circumvent the sign problem in QMC simulations of Fermi systems [46, 47, 48, 49, 50]. The fixed-node approximation is a popular approach and has yielded results for various systems [51, 52, 53, 54, 55, 56, 57, 42, 43], but introduces an intrinsic bias through an a priori chosen nodal structure.
This work establishes NN-VMC as a powerful tool for quantum simulations, significantly improving upon fixed-node DMC by refining nodal structures and reducing variational bias. Additionally, when the quantity of interest does not commute with the Hamiltonian, the DMC method requires the use of extrapolated estimators, which inherently introduce additional bias into the results. Our method achieves statistically superior ground-state energy estimates for QDs with up to 30 electrons, demonstrating the feasibility of machine learning in addressing key challenges in many-body physics.
Future research will focus on optimizing the computational efficiency of NN-VMC to scale to larger systems and further refine the learned nodal structures. This includes investigating nodal curvature metrics, improving algorithmic scalability, and leveraging GPU-accelerated frameworks such as PyTorch and TensorFlow to enhance performance. Additionally, NN-VMC can be extended to study spectral properties and excited states, broadening its applicability to strongly correlated fermionic systems.
Integrating machine learning with quantum simulations advances next-generation computational techniques, benefiting quantum computing and materials discovery.
Acknowledgements - SAV and WF acknowledge financial support from the Brazilian agency, Fundação de Amparo à Pesquisa do Estado de São Paulo grants #2023/07225-0 and #2020/10505-6, São Paulo Research Foundation (FAPESP). This work used Delta at the National Center for Supercomputing Applications through allocation CIS230072 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296 [58].